Project

General

Profile

Actions

RunPlan20150328 » History » Revision 6

« Previous | Revision 6/15 (diff) | Next »
Boyd, Steven, 22 March 2015 09:44


Target and Trigger timing + ToF calibration readiness - 28th/29th March 2015

The purpose of the run is to:

  • Get the system to a state where ToF (and CKOV) Calibration data can be taken
  • Take some TOF calibration data

Daily schedule

Date Saturday Saturday Sunday Sunday
estimated time 8am - 4pm 4 - 12pm 8am - 4pm 4pm - 12pm
MOM Y. Karadzhov Y. Karadzhov Y. Karadzhov Y. Karadzhov
Lead shifter P. Franchini P. Hanlet P. Frachini P. Hanlet
Shifter P. Soler P. Soler
BLOC
SOC
Expert Y Karadzhov (DAQ)
Aim Take Data
Beam? Yes Yes Yes Yes
Additional Personnel

Prerequisites

Run Preparation List

Data Requested

Beamline Configurations

  • N.B. Decay Solenoid will not be available.
  • We expect to be running with 2V beamloss and beam bump active.

Preparation

  • We will run beam into the hall so need to secure the MICE Hall
  • Target needs to be ready
  • BeamLine elements need to be ready
  • Detectors need to be ready
  • Controls needs to be ready
  • DAQ needs to be ready
  • OnMon and OnRec need to be ready
  • Data Mover needs to be ready

Shifter target training

  • at shift handover on Saturday - somewhere between 15:00-17:00

DAQ testing

  • The main goal of this run is the commissioning of the new trigger system, which requires a calibration of the Time-of-flight system..
  • We will take data with TOF0, TOF1 and Ckov. This means that we will use only one LDC (VME crate).
  1. We will start with a standalone readout of the TOF0/1 crate. This will be used for detecting eventual problems and debugging. The beam line will be set for Reference Run without Decay Solenoid.
  2. When we are satisfied with the quality of the data, we will switch to the original DAQ system and will take date for ~500 target pulses.

Reference Run without Decay Solenoid

Particle Species p at Tgt p@D1 p@D2 p@Tof0 Proton Absorber Q1 Q2 Q3 D1 DS D2 Q4 Q5 Q6 Q7 Q8 Q9 Similar Run(s)
MeV/c MeV/c MeV/c MeV/c mm A A A A A A A A A A A A DAQ run #
pion 272.42 268.67 265.98 143.4 29 57.8 105.6 64.6 202.8 --- 100.4 176.8 237.1 157.2 157.6 238.5 203.7 5729

TOF calibration

  1. The beam line will be set for TOF Calibration run
  2. The goal will be to take 30 000 target pulse or more.

TOF Calibration run

Number of pulses Beamloss Particle Species p at Tgt p@D1 p@D2 Proton Absorber Q1 Q2 Q3 D1 DS D2 Q4 Q5 Q6 Q7 Q8 Q9 Similar Run(s)
# V mm A A A A A A A A A A A A A A A A DAQ run #
30000 1 Positron 300.0 259.7 249.58 29 56.35 102.7 62.65 195.72 --- 98.59 163.16 218.79 114.14 83.40 125.70 80.06 1678

Finish

  • Shutdown target and re-open Hall for Monday

Updated by Boyd, Steven over 8 years ago · 6 revisions